Tuesday, Mar. 08: SWING Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans


  • Elastic data
  • Production data (Kin48_2)

Specific mission

      • Production run, Kin48_2
    • After Alex's work, Compton laser will be turned OFF !
  • Polarized beam at 10 uA on LH2 15 cm!
  • DVCS is running kin_48_2 [here]
  • LHRS is partially detuned: [here]
- LHRS=3.996GeV, Q1/Q2/Q3=1916.640(readout:1900)/1600/1600 (Q1 regul.OFF)
- PreScaleR : PS1=PS2=PS3=1, && PS8=200
- PreScaleL : PS9 =1, + DVCS config.
  • beam locks: no position lock, energy lock in Arc 2: [here]
  • DVCS daq configuration is [this one], we are measuring about 2% DT at 15uA.

BPM/Spot++ Check

  • After beam lost for considerable time (like beam study, pass change,...)
  • Ask MCC to run 5uA wth carbon hole target and make sure you can see the hole through Spot++

Please check more info on the white board in the Counting House ~!!